The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                    http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 42 March 12, 2000

Contents

News  1
Abstracts of 10 refereed papers 2
Conference contribution  8
Job opportunity  9

News

2MASS Second Incremental Data Release

The Two Micron All Sky Survey (2MASS) Second Incremental Data Release is now available for public access. This Release includes data from the northern and southern 2MASS facilities, covering 47% of the sky. The Release data products consist of a Point Source Catalog containing positions and photometry for over 162 million objects, an Extended Source Catalog containing positions, photometry and basic shape information for over 585,000 resolved sources, approximately 1.9 million compressed 512 × 1024 pixel (1′′/pixel) images in the three Survey bandpasses, and non-compressed “postagestamp” images for all of the Extended Source Catalog objects.

Most importantly, Image and Catalog data for large areas including and surrounding both Magellanic Clouds are included in this release.

The release data products can be accessed on-line from the IPAC/2MASS Web site at http://www.ipac.caltech.edu/2mass/ or directly from the NASA/Infrared Science Archive site at http://irsa.ipac.caltech.edu/. In the near future, the release Catalogs will be available via ftp download, and on a limited distribution DVD-ROM. Access to the 2MASS Atlas Images is currently possible only via the on-line services.

The 2MASS/IPAC webpage contains general information about this data release, including an on-line Explanatory Supplement (available at

http://www.ipac.caltech.edu/2mass/releases/second/doc/explsup.html), sky coverage maps,

images of interesting objects in the release, catalog characteristics, etc. A tool for determining whether a specified position is included in the release area is available on the NASA/Infrared Science Archive webpage. Questions about the release can be directed to the 2MASS Help Desk at

2mass@ipac.caltech.edu.

We encourage you to notify us (at 2mass@ipac.caltech.edu) about any refereed publications or conference proceedings (even in preprint form) which make use of these or earlier 2MASS Release data products. We will gladly provide links to your papers from the 2MASS website. Thank you very much in advance.

The 2MASS Project

The Two Micron All Sky Survey is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology. Funding for the survey has been provided by the National Aeronautics and Space Administration and the National Science Foundation.

Abstracts of Refereed Papers

Cold Atomic Gas in the Small Magellanic Cloud

John M. Dickey^{1} Ulrich Mebold^{2} Snezana Stanimirovic^{3} and Lister Staveley-Smith^{3}

1 Astronomy Department, University of Minnesota, 116 Church St. SE, Minneapolis, MN 55455, USA 2 Radioastronomisches Institut, Universität Bonn, Auf dem Hügel 71, D53121-Bonn, Germany 3 Australia Telescope National Facility, CSIRO, PO Box 76, Epping, NSW 2121, Australia

In this project we determine the mixture of thermal phases in the atomic medium of the Small Magellanic Cloud. The larger purpose is to study the difference in properties of the interstellar medium in a dwarf irregular galaxy compared with the Milky Way. The observations used the Australia Telescope Compact Array to measure absorption spectra of the λ21-cm line through the SMC, for comparison with λ21-cm emission spectra from a combination of data from the Parkes and Compact Array telescopes. Using the two together, we find the abundance of cool phase gas to be less than 15% of the total Hi. Thus the cool phase gas is very scarce, only about half as abundant relative to the total Hi as in the solar neighborhood, a lower figure than that found in the LMC, M31, M33, or in the solar neighborhood of the Milky Way. This is in qualitative agreement with the heating/cooling equilibrium calculation of Wolfire et al. (1995) for conditions in a dwarf irregular galaxy. In addition, we find the temperature of the cool phase clouds to be very low, typically 40 K or less. This is much less than the nominal diffuse cloud temperature in the Milky Way (50 to 100 K). This suggests that the Hi may be more abundant in dense, cold clouds in the SMC than in the Milky Way, where clouds of similar density and temperature are largely molecular.

Accepted by: The Astrophysical Journal

ROSAT HRI catalogue of X-ray sources in the LMC region

Manami Sasaki^{1}, Frank Haberl^{1}, and Wolfgang Pietsch^{1}

1 Max–Planck–Institut für extraterrestrische Physik, Giessenbachstraße, 85748 Garching, Germany

All 543 pointed observations of the ROSAT High Resolution Imager (HRI) with exposure times higher than 50 sec and performed between 1990 and 1998 in a field of 10 deg x 10 deg covering the Large Magellanic Cloud (LMC) were analyzed. A catalogue was produced containing 397 X-ray sources with their properties measured by the HRI. The list was cross-correlated with the ROSAT Position Sensitive Proportional Counter (PSPC) source catalogue presented by Haberl & Pietsch (1999) in order to obtain the hardness ratios for the X-ray sources detected by both instruments. 138 HRI sources are contained in the PSPC catalogue, 259 sources are new detections. The spatial resolution of the HRI was higher than that of the PSPC and the source position could be determined with errors mostly smaller than 15 arcsec which are dominated by systematic attitude errors. After cross-correlating the source catalogue with the SIMBAD data base and the TYCHO catalogue 94 HRI sources were identified with known objects based on their positional coincidence and X-ray properties. Whenever more accurate coordinates were given in catalogues or literature for identified sources, the X-ray coordinates were corrected and the systematic error of the X-ray position was reduced. For other sources observed simultaneously with an identified source the coordinates were improved as well. In total the X-ray position of 254 sources could be newly determined. The catalogue contains 39 foreground stars, 24 supernova remnants (SNRs), five supersoft sources, nine X-ray binaries, and nine AGN well known from literature. Another eight sources were identified with known candidates for these source classes. Additional 21 HRI sources are suggested in the present work as candidates for SNR, X-ray binary in the LMC, or background AGN because of their extent, hardness ratios, X-ray to optical flux ratio, or flux variability.

Accepted by: Astronomy & Astrophysics Supplements For preprints, contact manami@mpe.mpg.de

Broad-Band X-Ray Study of a Transient Pulsar RX J0059.2−7138

Makoto Kohno^{1}, Jun Yokogawa^{1} and Katsuji Koyama^{12}

1 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto 606-8502, Japan 2 CREST, Japan Science and Technology Corporation (JST), 4-1-8 Honmachi, Kawaguchi, Saitama 332-0012, Japan

We report on the results of the ASCA and ROSAT observations on RX J0059.2−7138, a transient X-ray pulsar in the Small Magellanic Cloud. The barycentric pulse period has been precisely determined to be 2.763221±0.000004 s. The pulse shape is almost identical in all of the energy bands. The pulse fraction increases with the photon energy below ∼ 2 keV, while it is nearly constant at ∼ 37% above ∼ 2 keV. The X-ray spectrum has been found to consist of two components. One is dominant above 2 keV, and exhibits sinusoidal pulsations. This component is well described by a typical model found in many X-ray binary pulsars, a power-law of photon index 0.4 with an exponential cut-off at 6.5 keV. The other is dominant below 1 keV and shows no significant pulsation. This component is represented by either a broken power-law with photon indices of 2.6 and 5.1 below and above a break energy of 0.9 keV, or a metal-poor thin-thermal plasma with a temperature of 0.37 keV. The phase-averaged luminosity is ∼ 10^{38} erg s^{−1} (0.1–10.0 keV) for both components. A hint of oxygen over-abundance is found in the absorbing column, possibly due to circumstellar gas ejected from an evolved companion.

Accepted by: Publications of the Astronomical Society of Japan

For preprints, contact kohno@cr.scphys.kyoto-u.ac.jp
Also available from the URL http://arXiv.org/abs/astro-ph/0003126
http://www-cr.scphys.kyoto-u.ac.jp/member/jun/job/

The Evolution of Rotating Stars

André Maeder and Georges Meynet

Geneva Observatory, CH–1290 Sauverny, Switzerland

First, we review the main physical effects to be considered in the building of evolutionary models of rotating stars on the Upper Main-Sequence (MS). The internal rotation law evolves as a result of contraction and expansion, meridional circulation, diffusion processes and mass loss. In turn, differential rotation and mixing exert a feedback on circulation and diffusion, so that a consistent treatment is necessary.

We review recent results on the evolution of internal rotation and the surface rotational velocities for stars on the Upper MS, for red giants, supergiants and W–R stars. A fast rotation is enhancing the mass loss by stellar winds and reciprocally high mass loss is removing a lot of angular momentum. The problem of the “break–up” or Ω–limit is critically examined in connection with the origin of Be and LBV stars. The effects of rotation on the tracks in the HR diagram, the lifetimes, the isochrones, the blue to red supergiant ratios, the formation of W–R stars, the chemical abundances in massive stars as well as in red giants and AGB stars, are reviewed in relation to recent observations for stars in the Galaxy and Magellanic Clouds. The effects of rotation on the final stages and on the chemical yields are examined, as well as the constraints placed by the periods of pulsars. On the whole, this review points out that stellar evolution is not only a function of mass M and metallicity Z, but of angular velocity Ω as well.

Accepted by: Annual Review of Astronomy & Astrophysics Vol. 38 (2000)

For preprints, contact Andre.Maeder@obs.unige.ch or Georges.Meynet@obs.unige.ch

Determination of masses and radii of the massive eclipsing binary system HV 2543

P.G. Ostrov^{1}, E. Lapasset^{2} and N.I. Morrell^{1}

1 Facultad de Ciencias Astronómicas y Geofísicas, Paseo del Bosque S/N, 1900 La Plata, Argentina 2 Observatorio Astronómico, Universidad Nacional de Córdoba, Laprida 854, 5000 Córdoba, Argentina

We present a V light curve of the eclipsing binary system HV 2543 (Sk–67^{∘}117) in the Large Magellanic Cloud based on CCD images acquired between the years 1995 and 1998. We have analysed this light curve and published radial velocity data, finding that this system is semidetached, with the secondary (less massive and less luminous component) filling its Roche lobe. From our analysis with the Wilson-Devinney code, we estimated the following masses and radii for the components of HV 2543: M_{1} = 25.63 ± 0.7 M_{⊙}, R_{1} = 15.54 ± 0.4 R_{⊙}, M_{2} = 15.63 ± 1.0 M_{⊙} and R_{2} = 13.99 ± 0.4 R_{⊙}. On the basis of B and V photometry of the field stars, we found that HV 2543 is member of an OB association, perhaps related to which the massive binary system Sk–67^{∘}105 belongs.

Accepted by: Astronomy & Astrophysics

For preprints, contact ostrov@fcaglp.edu.ar
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0003118

High Velocity Star Formation in the LMC

D. Graff^{1} and A. Gould^{1}

1 Ohio State University Departments Physics and Astronomy, Columbus, OH 43210, USA

Light-echo measurements show that SN1987A is 425 pc behind the LMC disk. It is continuing to move away from the disk at 18 km s^{−1}. Thus, it has been suggested that SN1987A was ejected from the LMC disk. However, SN1987A is a member of a star cluster, so this entire cluster would have to have been ejected from the disk. We show that the cluster was formed in the LMC disk, with a velocity perpendicular to the disk of about 50 km s^{−1}. Such high velocity formation of a star cluster is unusual, having no known counterpart in the Milky Way.

Submitted to: Astrophysical Journal Letters For preprints, contact graff.25@osu.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0002285

Is the [CO] index an age indicator for star forming galaxies?

L. Origlia^{1} and E. Oliva^{2}

1 Osservatorio Astronomico di Bologna, Via Ranzani 1, I–40127 Bologna, Italy 2 Osservatorio di Arcetri, Largo E. Fermi 5, I-50125 Firenze, Italy

The classical [CO] index, i.e. the strength of the Δv=2 CO absorption bands starting at 2.29 μm, is sometimes used to constrain the maximum age of star formation events in galaxies. In this paper we critically analyze

i) theoretical models which could predict either a factor of >2 drop or a pronounced increase of [CO] at ages older than 100 Myr, depending on the evolutionary tracks one adopts.

ii) observational data for young clusters in the LMC which do not show any strong relationship between the CO index and cluster age.

The above scenario indicates that the value of [CO] does not provide a reliable tool for estimating the age of stellar populations older than ∼10 Myr, i.e. after the first red supergiants have been formed. The contradictory results of theoretical models reflect problems in treating the evolution along the Asymptotic Giant Branch (AGB). In particular, those evolutionary synthesis models using stellar tracks which do not include the thermal pulsing AGB phase produce too weak CO features at 100–1000 Myr, i.e. in the range of ages when the near infrared emission is dominated by thermal pulsing AGB stars.

Accepted by: Astronomy & Astrophysics

For preprints, contact oliva@arcetri.astro.it
Also available from the URL http://www.arcetri.astro.it/ oliva

Multi-colour P L-relations of Cepheids in the hipparcos catalogue and the distance to the LMC

M.A.T. Groenewegen^{1} and R.D. Oudmaijer^{2,3}

1 Max-Planck Institut für Astrophysik, Karl-Schwarzschild-Straße 1, D-85740 Garching, Germany 2 Department of Physics and Astronomy, University of Leeds, LS2 9JT Leeds, U.K. 3 Blackett Laboratory, Imperial College of Science, Technology & Medicine, Prince Consort Road, London SW7 2BZ, U.K.

We analyse a sample of 236 Cepheids from the hipparcos catalog, using the method of “reduced parallaxes” in V, I, K and the reddening-free “Wesenheit-index”. We compare our sample to those considered by Feast & Catchpole (1997) and Lanoix et al. (1999), and argue that our sample is the most carefully selected one with respect to completeness, the flagging of overtone pulsators, and the removal of Cepheids that may influence the analyses for various reasons (double-mode Cepheids, unreliable hipparcos solutions, possible contaminated photometry due to binary companions).

From numerical simulations, and confirmed by the observed parallax distribution, we derive a (vertical) scale height of Cepheids of 70 pc, as expected for a population of 3-10 M_{⊙} stars. This has consequences for Malmquist- and Lutz-Kelker (Lutz & Kelker 1973, Oudmaijer et al. 1998) type corrections which are smaller for a disk population than for a spherical population.

The V and I data suggest that the slope of the Galactic P L-relations may be shallower than that observed for LMC Cepheids, either for the whole period range, or that there is a break at short periods (near log P_{0} ≈ 0.7 − 0.8).

We stress the importance of two systematic effects which influence the distance to the LMC: the slopes of the Galactic P L-relations and metallicity corrections. In order to assess the influence of these various effects, we present 27 distance moduli (DM) to the LMC. These are based on three different colours (V, I, K), three different slopes (the slope observed for Cepheids in the LMC, a shallower slope predicted from one set of theoretical models, and a steeper slope as derived for Galactic Cepheids from the surface-brightness technique), and three different metallicity corrections (no correction as predicted by one set of theoretical models, one implying larger DM as predicted by another set of theoretical models, and one implying shorter DM based on empirical evidence). We derive DM between 18.45 ± 0.18 and 18.86 ± 0.12. The DM based on K are shorter than those based on V and I and range from 18.45 ± 0.18 to 18.62 ± 0.19, but the DM in K could be systematically too low by about 0.1 magnitude because of a bias due to the fact that NIR photometry is available only for a limited number of stars.

From the Wesenheit-index we derive a DM of 18.60 ± 0.11, assuming the observed slope of LMC Cepheids and no metallicity correction, for want of more information.

The DM to the LMC based on the parallax data can be summarised as follows. Based on the P L-relation in V and I, and the Wesenheit-index, the DM is

18.60 ± 0.11 (±0.08 slope)(_{−0.15}^{+0.08} metallicity),

which is our current best estimate. Based on the P L-relation in K the DM is 18.52 ± 0.18 (±0.03 slope)(±0.06 metallicity)(_{−0}^{+0.10} sampling bias).

The random error is mostly due to the given accuracy of the hipparcos parallaxes and the number of Cepheids in the respective samples. The terms between parentheses indicate the possible systematic uncertainties due to the slope of the Galactic P L-relations, the metallicity corrections, and in the K-band, due to the limited number of stars. Recent work by Sandage et al. (1999) indicates that the effect of metallicity towards shorter distances may be smaller in V and I than indicated here.

From this, we point out the importance of obtaining NIR photometry for more (closeby) Cepheids, as for the moment NIR photometry is only available for 27% of the total sample. This would eliminate the possible bias due to the limited number of stars, and would reduce the random error estimate from 0.18 to about 0.10 mag. Furthermore, the sensitivity of the DM to reddening, metallicity correction and slope are smallest in the K-band.

Accepted by: Astronomy & Astrophysics

For preprints, contact groen@mpa-garching.mpg.de

Stellar Populations in the Large Magellanic Cloud from 2MASS

Sergei Nikolaev and Martin D. Weinberg

University of Massachusetts, Amherst, MA 01002, USA

We present a morphological analysis of the feature-rich 2MASS LMC color-magnitude diagram, identifying Galactic and LMC populations and estimating the density of LMC populations alone. We also present the projected spatial distributions of various stellar populations. The conditions that prevailed when 2MASS observed the LMC provided 10σ limiting sensitivity for J < ∼ 16.3, H < ∼ 15.3, K_{s} < ∼ 14.7. Major populations are identified based on matching morphological features of the colormagnitude diagram with expected positions of known populations, isochrone fits, and analysis of the projected spatial distributions.

2MASS has detected a significant population of asymptotic giant branch (AGB) stars (> ∼ 10^{4} sources) and obscured AGB stars (> ∼ 2 × 10^{3} sources). The LMC populations along the first-ascent red giant branch (RGB) and AGB are quantified. Comparison of the giant luminosity functions in the bar and the outer regions of the LMC shows that both luminosity functions appear consistent with each other. The luminosity function in central (bar) field has well-defined drop-off near K_{s} = 12.3^{m} corresponding to the location of RGB tip; the same feature is seen in the luminosity function of the entire LMC field.

Isochrone fits for the corresponding giant branches reveal no significant differences in metallicities and ages between central and outer regions of the LMC. This may be evidence for strong dynamical evolution in the last several gigayears. In particular, the observed LMC giant branch is well-fit by published tracks in the CIT/CTIO system with a distance modulus of μ = 18.5 ± 0.1, reddening E_{B−V} = 0.15 − 0.20, metallicity Z = 0.004_{−0.001}^{+0.002} and age 3 − 13 Gyr. Analysis of deep 2MASS engineering data with six times the standard exposure produces similar estimates.

Submitted to: The Astrophysical Journal

For preprints, contact nikolaev@redtail.astro.umass.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0003012

Magellanic Cloud Periphery Carbon Stars IV: The SMC

W.E. Kunkel^{1}, Serge Demers^{2} and M.J. Irwin ^{3}

1 Carnegie Institution of Washington, Chile 2 Université de Montréal, Canada 3 Institute of Astronomy, UK

The kinematics of 150 carbon stars observed at moderate dispersion on the periphery of the Small Magellanic Cloud are compared with the motions of neutral hydrogen and early type stars in the Inter-Cloud region. The distribution of radial velocities implies a configuration of these stars as a sheet inclined at 73±4 degrees to the plane of the sky. The near side, to the South, is dominated by a stellar component; to the North, the far side contains fewer carbon stars, and is dominated by the neutral gas. The upper velocity envelope of the stars is closely the same as that of the gas. This configuration is shown to be consistent with the known extension of the SMC along the line of sight, and is attributed to a tidally induced disruption of the SMC that originated in a close encounter with the LMC some 0.3 to 0.4 Gyr ago. The dearth of gas on the near side of the sheet is attributed to ablation processes akin to those inferred by Weiner & Williams (1996) to collisional excitation of the leading edges of Magellanic Stream clouds. Comparison with kinematic data of carbon stars, older than the LMC/SMC interaction, from Hardy, Suntzeff & Azzopardi (1989), and objects younger than this, such as supergiants of Maurice, Martin, & Bouchet (1989) and Cepheids by Mathewson et al (1986, 1988) leaves little doubt that forces other than gravity play a role in the dynamics of the H i.

Submitted to: The Astronomical Journal For preprints, contact Bill Kunkel kunkel@jeito.lco.cl

Abstracts of Non-Refereed Papers

The dependence of mass loss on the stellar parameters

Henny J.G.L.M. Lamers^{1}, Tiit Nugis^{2}, Jorick Vink^{1} and Alex de Koter^{3}

1 Astronomical Institute, University of Utrecht, Princetonplein5, NL-3584CC, Utrecht, The Netherlands 2 Tartu Observatory, 61602 Toravere, Estonia 3 Astronomical Institute, University of Amsterdam, Kruislaan 403, NL-1098SJ, Amsterdam, The Netherlands

We discuss the dependence of the mass-loss rates of OBA-stars, Central Stars of Planetary Nebulae and Wolf-Rayet stars on the stellar parameters. The winds of OBA stars have bi-stability jumps near T_{eff} ≃ 21 000 and 10 000 K, where the terminal velocity, and probably also the mass loss, changes drastically. This indicates a change of force multiplier parameters at these temperatures. We derive linear regression relations to describe the modified wind momentum and the terminal wind velocity of the OBA-stars in the Galaxy and we derive the metallicity dependence from the comparison with LMC and SMC stars. The modified wind momentum of the CSPN stars separates into two groups: one with a 30 times higher modified wind momentum than the other group. We derive regression relations for the momentum transfer efficiency η and the mass-loss rates of WR-stars, using clumpingcorrected mass-loss rates. Both η and Ṁ depend on L but also on composition, Y and Z. We discuss the possible reasons for the high mass-loss rates of WR-stars.

Accepted by: ”Thermal and Ionization Aspects of Flows from Hot Stars: Observations and Theory”, ASP Conf Series 204, H.J.G.L.M. Lamers & A. Sapar (eds), p. 305. For preprints, contact lamers@astro.uu.nl

Job Opportunity

Postdoctoral Research Associate at the University of Minnesota, Minneapolis

The University of Minnesota (UM) Astronomy Department, Minneapolis, MN seeks a Postdoctoral Research Associate in Observational Astronomy to begin in the Fall of 2000. The initial appointment will be for one year, with continued employment possible for at least two additional years. The primary responsibility of the individual will be to collaborate with Evan Skillman on space-based observations relevant to the understanding of the evolution of dwarf galaxies. Work with additional supporting observations from ground-based facilities will also be supported.

The applicant should have a Ph.D. in Physics, Astrophysics, or Astronomy, and completion of the Ph.D. degree is required before employment can begin. Preference will be given to applicants with a publication record in stellar populations or interstellar medium studies. Application materials, including three letters of recommendation, should be directed by 1 April 2000 to Evan D. Skillman, at the address below. The University of Minnesota is an equal opportunity educator and employer.

Dr. Evan D. Skillman

University of Minnesota

Department of Astronomy

116 Church Street, SE

Minneapolis, MN 55455, USA

   E-mail: skillman@astro.spa.umn.edu